Q 8: Assume that the population has a normal distribution. Estimate the population mean annual earning.

College students' annual earnings: 95% confidence; n=9,

sample mean = $21361, s=$865

Answer choices:

$20696 < u < $22026

$20796 < u < $21926

$19681 < u < $20128

$20584 < u < $21657
Answer #1

Solution :

degrees of freedom = n - 1 = 9 - 1 = 8

t\alpha/2,df = t0.025,8 = 2.306

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 2.306 * ( 865/ \sqrt 9)

Margin of error = E = 665

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

21361 - 665 < \mu < 21361 + 665

( $20696 < \mu < $22026 )